Find the section of the conicoid x^2/2-2y^2/3 = 2z by the plane x-2y+z = 1. What conic does this section represent? Justify your answer.
1
Expert's answer
2021-10-29T03:01:06-0400

"z=2y-x+1"

"x^2\/2-2y^2\/3=2(2y-x+1)"


"3x^2-4y^2=24y-12x+12"

"3(x^2+4x+4)-12-4(y^2+6y+9)+36=12"

"3(x+2)^2-4(y+3)^2=-12"


"\frac{(x+2)^2}{4}-\frac{(y+3)^2}{3}=-1"


This is a equation of conjugated hyperbola.

centre: "(-2,-3)"


semi-axes:

"a=2,b=\sqrt 3"